Name: Kevin Sutman

Content and Structure – The paper

1. Has a clearly stated topic or question that it will address – 9

2. Does address the topic or question – 5

3. Has a clearly articulated and logical structure:

a. First paragraph lays out the topic or question – 9 b. Later paragraphs have topic sentences and clearly identifiable main ideas – 9 c. Transitions between paragraphs are logical – 9

4. Accurately presents facts and viewpoints of sources – 10

5. Provides sufficient textual evidence to support the argument – 10

6. Is concise, clear, and not too repetitive – 10

7. Has a clearly identifiable conclusion – 10

8. Has a convincing argument that draws together the evidence presented – 13

Total: 94 (A)

Style and English Usage – The paper

1. Is clear, direct, and appropriate in tone – 10

2. Is free of misspellings – 10

3. Is free of grammatical mistakes – 10

4. Avoids incomplete sentences – 10

5. Avoids run-on or overly complicated sentences - 9

6. Uses correct punctuation – 9

7. Is free of jargon, cliches, and misused words – 14

8. Cites quotations and/or references correctly – 9

Total: 96 (A)

Comments: Kevin, I’m still dubious about the “who told you you could sell my spleen?” as a direct quotation. It is the title of that chapter, but Skloot is not claiming that it is a direct quotation. Your ultimate point, that the medical treatment Henrietta received was good, but that the unauthorized use of her cells was not ethical, seems reasonable to me. It would have been better to lay out exactly that distinction in the first paragraph, though. That would have set up exactly the point you wanted to make with the essay more clearly. This is a good improvement over the first version, though.